robyn is going skiing. A T-Bar is used to pull the skiers up the hill. the bill has an incline of 10.0 degrees above the horizontal. she starts from rest and has a mass of 50.0kg. the coefficient of friction between her skis and the snow is 0.100. the force on the cable of the T-bar is applied at 30.0 degrees to the surface of the hill. if the T-bar accelerates her at 3.92m/s2, find the force applied to the cable.


1
Expert's answer
2021-10-31T18:12:33-0400

Apply Newton's second law:


"ma=-f-mg\sin10°+F\cos(30°-10°),\\
f=\mu N=\mu mg\cos10°.\\\space\\
F=\frac{m(a+g\sin10°+\mu g\cos10°)}{\cos20°}=350\text{ N}."